OpenStudy (anonymous):

solve for x by factoring. (x+2)(x+8)=0. show your work please :)

OpenStudy (anonymous):

x=-2 and -8. I got this by setting x = 2 and x=8 to 0 and then solving for x.

OpenStudy (anonymous):

FOIL first dawggy! so you get x^2 plus 10x plus 16

OpenStudy (accessdenied):

Your function is already factored completely... We can use the 'zero-product property' to solve for x. Zero-product just states that if a*b = 0, then a= 0 or b=0 x + 8 = 0 or x + 2 = 0. x=-8 or x=-2.

OpenStudy (anonymous):

i was just about to say that. ^ lol . your already done factoring,, bruuh